OpenStudy (alexa21):

Solve the separable differential equation: 9x-4y sqrt(x^2+1) dy/dx = 0 Subject to the initial condition: y(0)=3.

OpenStudy (alexa21):

\[9x-4y \sqrt{x^2+1}\frac{ dy }{dx}=0\]

OpenStudy (irishboy123):

\(9x = 4y \sqrt{x^2+1}\frac{ dy }{dx}\) \(\dfrac{9x}{ \sqrt{x^2+1} } = 4y \frac{ dy }{dx}\) then do !

OpenStudy (irishboy123):

pattern on LHS is \(\dfrac{d}{dx} \sqrt {f(x)} = \dots\)

OpenStudy (alexa21):

Thank you, but what do you mean "pattern on LHS"?

OpenStudy (irishboy123):

\[\implies \int \left ( \dfrac{9x}{ \sqrt{x^2+1} } = 4y \frac{ dy }{dx} \right) \color{red}{dx}\] @SolomonZelman best guy/gal at explaining this, IMHO :-))

OpenStudy (solomonzelman):

Yes, the equation is separated as follows (as IrishBoy123 showed). \(\color{black}{\displaystyle \frac{9x}{\sqrt{x^2+1}}=4y \frac{ dy }{dx}}\) then, you integrate both sides wrt x, \(\color{black}{\displaystyle \color{red}{\int}\frac{9x}{\sqrt{x^2+1}}\color{red}{dx}=\color{red}{\int}4y \frac{ dy }{dx}}\color{red}{dx}\) \(\color{black}{\displaystyle \int\frac{9x}{\sqrt{x^2+1}}dx=\int 4y~dy}\) -------------------------------------------------------------- (This is why, btw, when you solve equations in the form \(\color{black}{\displaystyle f(x)=g(y)~\frac{dy}{dx} }\) you integrate wrt x and wrt y on each side, as follows, \(\color{black}{\displaystyle \int f(x)~dx=\int g(y)~dy}\) then in general you would continue... \(\color{black}{\displaystyle F(x)+C=G(y)}\) (capital lettered functions are antiderivaties) The use your initial condition to solve for C ... etc. --------------------------------------------------------------- The integral on the left side is by u-sub (you can do trig-sub as well, as a matter of a fact), and the right is obvious. The integral on the right is by power rule. Then, use your initial conditions to find C.

OpenStudy (solomonzelman):

Nothing to add, other than I'm not the best source for DE material on this site ... hope this helps.

OpenStudy (irishboy123):

In terms of *Pattern*, via the power and chain rules.... \(\dfrac{d}{dx} (x^2 + 1)^{1/2} = \dfrac{1}{2} \cdot (x^2 + 1)^{\color{red}{-1/2}} \cdot 2x= \dfrac{x }{\sqrt{x^2 + 1}}\) So \(\int \dfrac{9x }{\sqrt{x^2 + 1}} ~ dx =9 \int \dfrac{x }{\sqrt{x^2 + 1}} ~ dx =9 \int \dfrac{d}{dx} (x^2 + 1)^{1/2} ~ dx\) \( =9 (x^2 + 1)^{1/2} + C\) It looks worse when it's written out longhand. But, in the head, it's way shorter. You can achieve the same result in the manner of a **substitution**: \(z = x^2 + 1, \color{blue}{dz} = 2x ~ \color{blue}{dx}\) that often includes treating the components of the Liebnitz differential notation as components of a fraction, so... \(9x-4y \sqrt{x^2+1} \dfrac{dy}{dx} = 0\) \(\implies 9x=4y \sqrt{x^2+1} \color{red}{\dfrac{dy}{dx}}\) \(\implies \color{red}{dx} ~ 9x=4y \sqrt{x^2+1} ~ \color{red}{dy}\) I'm no expert but these ideas are waved around so much; and I wonder why.